# Hierarchical Latent Class Models for Mortality Surveillance Using Partially Verified Verbal Autopsies

### v1

@@ Abstract
Monitoring data on causes of death is an important part of understanding the burden of diseases and effects of public health interventions. Verbal autopsy (VA) is a well-established method for gathering information about deaths outside of hospitals by conducting an interview to family members or caregivers of a deceased person. Existing cause-of-death assignment algorithms using VA data require either domain knowledge about the symptom-cause relationship, or large training datasets. When a new disease emerges, however, only limited information on symptom-cause relationship exists and training data are usually lacking, making it challenging to evaluate the impact of the disease. In this paper, we propose a novel Bayesian framework to estimate the fraction of deaths due to an emerging disease using VAs collected with partially verified cause of death. We use a latent class model to capture the distribution of symptoms and their dependence in a parsimonious way. We discuss potential sources of bias that may occur due to the cause-of-death verification process and adapt our framework to account for the verification mechanism. We also develop structured priors to improve prevalence estimation for sub-populations. We demonstrate the performance of our model using a mortality surveillance dataset that includes suspected COVID-19 related deaths in Brazil in 2021.
Keywords: Verification bias; structured prior; domain adaptation; semi-supervised learning.
@@ 1 Introduction
Monitoring data describing cause of death is an essential component for understanding the burden of disease and evaluating public health interventions. Only about two-thirds of deaths worldwide are registered and up to half of these deaths do not have an assigned cause, with the least information available from countries with the most need (World Health Organization,, 2021). A widely used tool to obtain information on causes of death when a medically certified cause-of-death is not available is verbal autopsies (VA). VA involves a structured questionnaire administered to family members or caregivers of a recently deceased person. The VA interviews collect information about the circumstances, signs, and symptoms leading up to death. VAs are widely adopted both in research settings and by national statistical offices (e.g., Maher et al.,, 2010; Nkengasong et al.,, 2020). Such information is either analyzed by physicians or with statistical algorithms to assign causes of death.
@@
During public health emergencies when a new disease emerges, VA is usually the only feasible tool to gather information on causes of death in many low-resource settings, e.g., during the Ebola hemorrhagic fever outbreaks (Alert et al.,, 2003), the Dengue epidemic (Saqib et al.,, 2014), and the COVID-19 pandemic (de Souza et al.,, 2020; Rosen et al.,, 2021). In such settings, physician review of VAs is usually infeasible since the process is time-consuming and expensive, especially during a public health crisis, and cause-of-death assignment needs to be performed by automatic VA algorithms in any medium- to large-scale mortality surveillance settings.
@@
While significant advances have been made in automatic cause-of-death assignment algorithms in the last several decades (Chandramohan et al.,, 2021), using VAs for the purpose of monitoring mortality patterns and trends due to a new disease remains challenging. First, all existing VA algorithms operate with a pre-defined set of symptoms and causes and their relationships are either provided by physicians and assumed to be known (Byass et al.,, 2019; McCormick et al.,, 2016) or need to be estimated from high-quality training data, i.e., reference deaths with known causes verified through a separate mechanism (Kunihama et al.,, 2020; Moran et al.,, 2021). In the case of a new disease, the relationship between symptoms and the new cause of death is generally unknown and needs to be estimated from the limited training data, i.e., reference deaths with verified cause of death, during the outbreak. Moreover, the subset of deaths with the cause of death verified may not be a random sample of the target population, due to logistical reasons and changing public health priorities. Such reference deaths can exhibit different symptom-cause relationships and VA algorithms trained on these deaths could lead to biased estimates if the selective sampling process is not properly accounted for. This issue is similar to the verification bias in the diagnostic test literature, where the accuracy of a diagnostic test can be distorted by evaluation based on the patients with verified disease status only (Zhou,, 1998). To our knowledge, there are no guidelines or recommendations in the literature to account for the selection process of reference death in VA research.
@@
Second, the target of inference for mortality surveillance is usually the time-varying prevalence of the disease, potentially further stratified by key demographic groups. Existing literature on VA analysis, however, typically focuses on estimating the cause-specific mortality fraction (CSMF) for a single target population with a stable distribution of symptoms and causes of death. During the course of an outbreak, the population mortality profiles can change rapidly, due to both disease dynamics and changing treatment and intervention measures. Thus the accuracy of cause-of-death assignment methods trained on data from the early phase of an outbreak may deteriorate over time. Therefore, it is important to flexibly adapt the model to the changing distribution over time and across related sub-populations. A related line of work by Moran et al., (2021) and Kunihama et al., (2024) shows that modeling covariate-dependent symptom distributions can improve VA algorithms in general. However, their target of inference is still the overall prevalence of a static population, instead of time-varying or sub-population prevalence. As we demonstrate in both simulation and real data analysis, directly extending single-population models by treating the sub-populations as independent can lead to estimates with large variations, especially when the sample size is small.
@@
Both challenges are closely related to the generalizability of VA algorithms from a non-representative training dataset to one or multiple target populations. In this paper, we address these two issues in a unified framework by extending the models in Li et al., (2024) and Wu et al., (2024) to a hierarchical framework to deal with the distribution shift across multiple sub-populations. We focus on estimating the prevalence of a single disease over time for sub-populations defined by key demographic indicators. We propose a hierarchical latent class model informed by the causal structure of the cause verification process to account for the heterogeneity of the collected data stream. Furthermore, we propose a novel set of structured priors on the sub-population prevalence to efficiently borrow information and improve the precision of estimates.
@@
The rest of the paper is organized as follows. Section 2 introduces our proposed nested latent class model with structured priors over sub-populations. Section 3 develops an efficient Markov Chain Monte Carlo (MCMC) algorithm for posterior sampling. Section 4 demonstrates the proposed model with a series of simulation experiments and a COVID-19 surveillance dataset from Brazil. Section 5 concludes with a discussion of limitations and future work on using VA for mortality surveillance.
2 Models
2.1 Verification mechanism
Let $X_{i}\in\{0,1\}^{p}$ denote the $p$-dimensional vector of signs/symptoms for the $i$-th death and $Y_{i}\in\{0,1\}$ denote the corresponding cause of death. The overall prevalence of the disease is $p(Y)$. We consider the case where only a subset of deaths were verified, i.e., assigned a cause by an independent mechanism. Let $L_{i}\in\{0,1\}$ be the indicator of whether the cause of death for the $i$-th death is known. We refer to the process generating $L$ as the verification mechanism thereafter. Existing literature typically assumes that $L_{i}=1$ for all deaths with VAs in some training populations, usually different from the target population where cause-of-death assignment is performed, and various assumptions between the training and target distributions are made to ensure transportability of the classifier (Li et al.,, 2024). Here we focus on a different scenario where a subset of the deaths in the target population have a known cause-of-death label $L$.
@@ 2.1 Verification mechanism
When the verified deaths, i.e., the training data, are a subset of the target population, it becomes crucial to understand how the deaths are selected for verification. If the training subset is not a simple random sample of the population, models estimated using the verified deaths may be inadequate to generalize to the entire population directly. In surveillance settings, it is common to over-sample deaths of certain characteristics, e.g., deaths in certain age groups or from certain locations, into the verification dataset in order to increase the chance of identifying certain deaths, and the selection process may also change over time. Valid inference is still possible if the verification mechanism is conditionally ignorable, i.e., it depends only on the observed quantities. Assumption 1 formally characterizes such scenarios.
@@
Assumption 1 is equivalent to assuming that the cause of death is missing at random (Rubin,, 1976). The choice of $D$ should be informed by how the reference deaths with verified causes are selected. When selection probability depends on known and observable variables, we can let $D$ be the strata defined by these variables. When the verification mechanism is not clearly specified, the choice of $D$ is contextual. Including more covariates in $D$ mitigates the bias from potential violation of the conditional ignorability assumption but it increases model complexity. In practice, proxy variables such as time period and region indicators may be included in $D$ to mitigate the effect from any unobserved variables that vary over time or space.
2.2 Semi-supervised approach
Given a choice of $D$ that renders the selection mechanism conditionally ignorable, it suffices to learn a binary classification rule using the training dataset. The population prevalence can be estimated by $\hat{p}(Y)=\frac{1}{n}\sum_{i=1}^{n}\hat{p}(Y\mid X=x_{i},D=d_{i},L=1)$. Sub-population prevalence $p(Y\mid D=d)$ can also be obtained by aggregation in a similar fashion. A major drawback of such discriminant classification models, however, is that they do not take into account the potentially large collection of unlabelled VA data. On the other hand, semi-supervised approaches that model both the labelled and unlabelled data can achieve better predictive performance in tasks where $Y$ is the ‘cause’ of the observed covariates $X$, as the distribution of $X$ in the unlabelled data also contains information about the conditional distribution of $X$ given $Y$ (Kügelgen et al.,, 2020). Classification tasks with this type of structure are usually referred to as being ‘anticausal’ (Schölkopf et al.,, 2012). The VA problem is anticausal in nature because the majority of the symptoms and indicators are consequences of the underlying cause of death (King and Lu,, 2008; Li et al.,, 2024). Figure 1(a) illustrates the data generating process of the observable quantities under the assumed anticausal structure and the conditional independence relationship in Assumption 1. Therefore, in this work, we treat the unknown cause of death $Y$ as missing data and fit our models on all the deaths together.
@@ 2.2 Semi-supervised approach
It is also worth noting that the targets of inference may depend on the stratification of the population by variables collected by the VA survey itself, such as sex and age of the deceased. Strictly speaking, the graphical model in Figure 1(a) can be refined by further partitioning the variables collected by VA into $X=(X_{C},X_{E})$ where $X_{C}$ are variables that affect the risk of the cause of death, e.g., demographic variables, and $X_{E}$ are the variables affected by the cause of death, e.g., medical symptoms. This leads to the data generating process shown in Figure 1(b). Previous work on VA models typically do not differentiate $X_{C}$ from $X_{E}$. The only two exceptions in the literature are Moran et al., (2021) and Kunihama et al., (2024). Both papers consider separating some covariates from the rest of the symptoms, and model the conditional distribution of covariates given cause of death. They have shown that explicit estimation of symptom distribution given covariates improves the estimation of the prevalence under latent Gaussian factor models. They aim to flexibly characterize the joint distribution of symptoms. Our motivation is different. We choose $X_{C}$ to be the pre-disease variables that define meaningful strata in the population, and our target of inference is the sub-population prevalence $p(Y\mid D,X_{C})$ instead of $p(Y)$. Therefore, instead of modeling the conditional distribution of covariates given cause of death as in previous work, we explicitly model the conditional distribution of cause of death given covariates $X_{C}$. This factorization also enables the use of more interpretable and structured prior specification for $p(Y\mid D,X_{C})$ described in Section 2.4. To simplify notation in the rest of the paper, we omit $X_{C}$ and $X_{E}$, and use $D$ to denote any variables defining stratification regardless of whether they are collected in the VA survey, and $X$ to denote the rest of the indicators from the VA survey.
2.3 Nested latent class model
We now turn to model specification. Let $D_{i}\in\{1,...,G\}$ denote the combination of all levels of the stratification variables. The anticausal structure in Figure 1 leads to the factorization of the joint distribution $p(X,Y,D)=p(D)p(Y\mid D)p(X\mid Y,D)$. Many existing VA models simplify the last conditional probability into $p(X\mid Y)$ by assuming the conditional distribution of symptoms are transferable (e.g., McCormick et al.,, 2016; Byass et al.,, 2019). Such assumptions are usually violated when considering symptom distributions across deaths of different sex and age groups (Moran et al.,, 2021) or over different populations (Li et al.,, 2024). To illustrate the conditional dependence on $D$, we use the COVID-19 dataset described in Section 4.2 to compute the empirical symptom prevalence and the Matthews correlation coefficient (Matthews,, 1975) of some selected symptoms across different age groups and over time. Figure 2 shows that the symptom prevalence varies across both time and age groups. For example, among COVID-19 related deaths, the proportion of observations with fever as a symptom decreases over time and as age increases. Figure 3 shows that the dependence among symptoms also varies significantly with the age and time period of the death.
@@ 2.3 Nested latent class model
Figure 3: Matthews correlation coefficient (Matthews,, 1975) of a subset of symptoms among deaths related to COVID-19 among deaths in three age groups from three time periods in the Brazil VA dataset. Missing data are removed in calculating correlation coefficient.
In order to capture the heterogeneity of symptoms distributions over different sub-populations in a parsimonious way, we adopt the nested latent class model framework proposed in Li et al., (2024) and Wu et al., (2024), where individual-level latent class membership indicator $Z_{i}\in\{1,2,...,K\}$ is introduced to flexibly capture the dependence of $X$. We assume the following data generating process,
@@
The introduction of the latent class nested under each cause of death leads to a flexible characterization of cause-specific dependent symptom profiles that vary across strata, i.e., after integrating out the latent indicators, the conditional distribution of symptoms given a cause of death and sub-population is $p(X_{i}\mid Y_{i}=c,D_{i}=g)=\sum_{k}\lambda_{ck}^{(g)}\prod_{j}\phi_{ckj}^{x_{ij}}(1-\phi_{ckj})^{1-x_{ij}}$. When $K=1$, the model reduces to the commonly used conditional independent model in McCormick et al., (2016). In order to facilitate easier interpretations and mitigate the risk of overfitting, we fix the latent response probabilities $\bm{\phi}$ to be invariant across strata and let the sub-populations mix differently over the latent symptom profiles. With sufficiently many latent classes $K$, this representation is flexible enough to capture any multivariate discrete distribution (Dunson and Xing,, 2009).
The prior specification of $\pi^{(g)}$ needs to be designed for the specific choice of stratification variable $D$. We discuss our structured prior for $\pi^{(g)}$ in the next subsection. For the rest of the latent parameters, we put stick-breaking priors on $\bm{\lambda}_{c}^{(g)}$ and conjugate Beta priors on $\bm{\phi}$, i.e.,

### v3

@@ Abstract
Monitoring cause-of-death data is an important part of understanding disease burdens and effects of public health interventions. Verbal autopsy (VA) is a well-established method for gathering information about deaths outside of hospitals by conducting an interview to caregivers of a deceased person. It is usually the only tool for cause-of-death surveillance in low-resource settings. A critical limitation with current practices of VA analysis is that all algorithms require either domain knowledge about symptom-cause relationships or large labeled datasets for model training. Therefore, they cannot be easily adopted during public health emergencies when new diseases emerge with rapidly evolving epidemiological patterns. In this paper, we consider estimating the fraction of deaths due to an emerging disease. We develop a novel Bayesian framework using hierarchical latent class models to account for the informative cause-of-death verification process. Our model flexibly captures the joint distribution of symptoms and how they change over time in different sub-populations. We also propose structured priors to improve the precision of the cause-specific mortality estimates for small sub-populations. Our model is motivated by mortality surveillance of COVID-19 related deaths in low-resource settings. We apply our method to a dataset that includes suspected COVID-19 related deaths in Brazil in 2021.
Keywords: Verification bias; structured prior; domain adaptation; quantification learning.
@@ 1 Introduction
Monitoring data describing cause of death is an essential component for understanding the burden of disease and evaluating public health interventions. Only about two-thirds of deaths worldwide are registered, and up to half of these deaths do not have an assigned cause, with the least information available from countries with the most need (World Health Organization,, 2021). A widely used tool to obtain information on causes of death when a medically certified cause-of-death is not available is verbal autopsies (VA). VA involves a structured questionnaire administered to family members or caregivers of a recently deceased person. The VA interviews collect information about the circumstances, signs, and symptoms leading up to death. VAs are widely adopted both in research settings and by national statistical offices (e.g., Maher et al.,, 2010; Nkengasong et al.,, 2020). Such information is either analyzed by physicians or with statistical algorithms to assign causes of death.
@@
During public health emergencies when a new disease emerges, VA is usually the only feasible tool to gather information on causes of death in many low-resource settings, e.g., during the Ebola hemorrhagic fever outbreaks (Alert et al.,, 2003), the Dengue epidemic (Saqib et al.,, 2014), and the COVID-19 pandemic (de Souza et al.,, 2020; Rosen et al.,, 2021). In such settings, physician review of VAs is usually infeasible since the process is time-consuming and expensive, especially during a public health crisis, and cause-of-death assignment needs to be performed by automatic VA algorithms in any medium- to large-scale mortality surveillance settings. While significant advances have been made in automatic cause-of-death assignment algorithms in the last several decades (see e.g., Chandramohan et al.,, 2021, for a review), using VAs for the purpose of monitoring mortality patterns and trends due to new diseases remains challenging and largely unexplored in the existing literature. The lack of guidance on using algorithms to analyze VA for rapid mortality surveillance presents a critical gap for governments and public health researchers, as evidenced by the latest COVID-19 pandemic.
@@
In order to be prepared for future outbreaks of new diseases, two critical challenges need to be addressed with VA algorithms. First of all, all existing VA algorithms operate with a pre-defined set of symptoms and causes and their relationships are either provided by physicians and assumed to be known (Byass et al.,, 2019; McCormick et al.,, 2016) or need to be estimated from high-quality training data, i.e., reference deaths with known causes verified through a separate mechanism (Kunihama et al.,, 2020; Moran et al.,, 2021). In the case of an emerging disease, the relationship between symptoms and the new cause of death is generally unknown and needs to be estimated from the limited training data during the outbreak. Unlike traditional settings, the reference deaths with verified causes during an outbreak are usually not a random sample of the population due to logistical constraints and changing public health priorities. Such reference deaths can exhibit different symptom-cause relationships, and VA algorithms trained on these deaths could lead to biased estimates. This issue is similar to the verification bias in the diagnostic test literature, where the accuracy of a diagnostic test can be distorted by evaluation based on the patients with verified disease status only (Zhou,, 1998). No guidelines exist on when and how to account for the selection process of reference death in VA research.
@@
In addition, unlike routine analysis of VA with a single and static population, it is often of interest to monitor how the fraction of deaths due to each disease changes over time and how it varies across fine-scale demographic groups. During the course of an outbreak, the population mortality profiles can change rapidly due to both disease dynamics and changing treatment and intervention measures. Thus, the accuracy of cause-of-death assignment methods trained on data from the early phase of an outbreak may deteriorate over time. It is important to flexibly adapt the model to the changing distribution over time and across related sub-populations. A related line of work by Moran et al., (2021) and Kunihama et al., (2024) shows that modeling covariate-dependent symptom distributions can improve VA algorithms in general. However, their target of inference is still the overall cause-specific mortality fraction in a static population, and covariates are introduced only to improve the estimation of the joint distribution of symptoms. Directly extending single-population models by treating the sub-populations as independent can lead to inefficient estimates when sub-population sample sizes are small.
@@
In this paper, we address these challenges through the lens of distribution shift. We show that the issue with emerging disease monitoring using VAs can be viewed as a lack of generalizability of algorithms from a non-representative training dataset to target populations (Li et al.,, 2024; Wu et al.,, 2024). We develop a unified hierarchical latent class modeling framework informed by the causal structure of the cause-of-death verification process. Our model allows direct estimation of the fraction of deaths due to the disease across multiple sub-populations over time. We also propose a novel set of structured priors to efficiently borrow information and improve the precision of estimates.
@@
The rest of the paper is organized as follows. Section 2 introduces a COVID-19 surveillance dataset from Brazil that motivates this study. Section 3 introduces our proposed hierarchical latent class model with structured priors over sub-populations. Section 4 develops an efficient Markov Chain Monte Carlo (MCMC) algorithm for posterior sampling. Section 5 demonstrates the proposed model with a series of analyses using both simulated data and the COVID-19 surveillance dataset from Brazil. Section 6 concludes with a discussion of limitations and future work on using VA for mortality surveillance.
2 Brazil COVID-19 surveillance data
Our study is motivated by the task of identifying COVID-19 deaths in a flu syndrome surveillance dataset collected in Brazil from January to October 2021. All deaths in this database had severe acute respiratory syndrome by COVID, other respiratory diseases or chronic diseases. The dataset contains the independently assigned underlying cause of death for $411,491$ deaths. The dataset consists of $p=14$ binary indicators for each death, including symptoms such as fever, vomiting, loss of taste and smell, etc. Additional information on vaccine, PCR and antigen test results are also recorded but they are largely missing and thus are excluded from our analysis. Within this population of deaths suspected of COVID-19, the proportion of deaths with the final classification of being COVID-19 related increases from January to April and declines afterward. We consider eight age groups, where the first includes deaths between 0 and 30 years old and the subsequent groups are defined by ten-year increments. We let the first age group span 30 years because deaths under $20$ years old are rare in this population. Figure 1 shows the true fraction of deaths due to COVID-19 among these suspected deaths, within each sub-population. Figure 2 shows that the sample sizes across sub-populations are notably imbalanced, with particularly small samples observed in January and October, as well as in the first and last age groups.
Figure 1: Fraction of deaths due to COVID-19 among the deaths suspected of COVID-19 across different age groups and sexes over time.
3 Methods
3.1 Assumptions on verification mechanism
Let $X_{i}\in\{0,1\}^{p}$ denote the $p$-dimensional vector of signs/symptoms for the $i$-th death and $Y_{i}\in\{0,1\}$ denote whether the underlying cause of death is COVID-19 related. The overall fraction of deaths due to COVID-19, or the cause-specific mortality fraction (CSMF) of COVID-19, is then $p(Y)$. In this paper, we focus on symptom surveys designed specifically to identify COVID-19 deaths. As a result, there is limited information available to distinguish among non-COVID-19 causes. Therefore, we treat $Y$ as a binary variable.
We focus on the situations where the cause of death was known for only a subset of deaths. Typically, these causes are assigned and verified by physicians based on extensive reviews of all available information for each death. Let $L_{i}\in\{0,1\}$ be the indicator of whether the cause of death for the $i$-th death is known. Existing literature typically assumes that $L_{i}=1$ for all deaths with VAs in some training dataset and $L_{i}=0$ for all deaths in the target dataset. When these two datasets arise from different populations, modeling assumptions on the two data distributions are necessary in order to ensure transportability of the classifier built on the training data (Li et al.,, 2024).
@@ 3.1 Assumptions on verification mechanism
In the context of mortality monitoring during outbreaks, the deaths with verified causes are usually a subset of the target population. It is crucial to understand how the deaths are selected into the training dataset. If the training subset is not a simple random sample of the population, models estimated using the verified deaths may be inadequate to generalize to the entire population directly. In surveillance settings, it is common to over-sample deaths of certain characteristics, e.g., deaths in certain age groups or from certain locations, in order to increase the chance of identifying certain types of deaths, or simply due to logistical reasons. The selection process may also change over time. We refer to any such process that generates $L$ as the verification mechanism.
Valid inference of the CSMF is possible if the verification mechanism is conditionally ignorable. Assumption 1 formally characterizes this condition.
@@
Assumption 1 is equivalent to assuming that the cause of death is missing at random (Rubin,, 1976). The choice of $D$ should be informed by how the reference deaths with verified causes are selected. In many situations with mortality surveillance, the selection mechanism is known and depends on observable quantities. When it is not the case, the choice of $D$ is contextual. Including more covariates in $D$ mitigates the bias from potential violation of the assumption but increases model complexity. Proxy variables such as time periods and region indicators may be included in $D$ to mitigate the effect due to unobserved variables that vary over time or space.
Figure 3(a) illustrates the data generating process of random variables under Assumption 1, where no arrows point from $Y$ to $L$ directly and there are no unobserved confounders. It is worth noting that in the data generating process, we assume the observed variables $X$ are generated conditional on the underlying cause of death $Y$, thus the arrow from $Y$ to $X$. This assumption follows the convention in VA analysis where symptoms are treated as consequences of the underlying disease (King and Lu,, 2008; Li et al.,, 2024). Prediction tasks with this type of structure are usually referred to as being anticausal (Schölkopf et al.,, 2012).
The anticausal structure can be further refined in the context of VA. Conceptually, we may partition the variables collected from VA into $X=(X_{C},X_{E})$ where $X_{C}$ are variables that affect the risk of the cause of death, e.g., demographic variables, and $X_{E}$ are the variables affected by the cause of death, e.g., medical symptoms. This leads to the data generating process with a mixed causal structure shown in Figure 3(b). Our target of inference in this setting is the sub-population CSMFs, $p(Y\mid D,X_{C})$, instead of the overall CSMF, $p(Y)$. Previous VA models have also explored separating demographic variables from medical symptoms, and such separation of covariates and symptoms has been shown to improve cause-of-death classification (Moran et al.,, 2021; Kunihama et al.,, 2024). However, their focus was on improving the estimation of the overall CSMF, and the covariates are still treated as downstream consequences of the cause of death in the assumed model structures. In contrast, our model explicitly distinguishes between upstream covariates and downstream symptoms, and explicitly targets covariate-dependent CSMFs. To simplify notation in the rest of the paper, we note that $X_{C}$ and $D$ play the same role in terms of the model, and differ only in whether they are collected from the VA survey or supplied as external information. Thus we omit $X_{C}$ and $X_{E}$, and use $D\in{1,...,G}$ to denote the combination of all levels of variables defining stratification, and $X$ to denote the rest of the indicators from the VA survey, as in Figure 3(a).
@@
3.2 Hierarchical latent class model
We now turn to model specification. Given a choice of $D$ that renders the selection mechanism conditionally ignorable, one may learn a binary classification rule using the training dataset. The overall CSMF can be estimated by $\hat{p}(Y)=\frac{1}{n}\sum_{i=1}^{n}\hat{p}(Y\mid X=x_{i},D=d_{i},L=1)$. Sub-population CSMFs $p(Y\mid D=d)$ can also be obtained by aggregation in a similar fashion. A major drawback of such discriminant classification models is that they do not take into account the large collection of unlabeled VA data, which provides rich information on the distribution of $X$ in anticausal problems (Kügelgen et al.,, 2020). Therefore, we consider a semi-supervised approach that models both labeled and unlabeled data together.
Our assumed data generating process in Figure 3 leads to the factorization of the joint distribution $p(X,Y,D)=p(D)p(Y\mid D)p(X\mid Y,D)$. Many existing VA models simplify the last conditional probability into $p(X\mid Y)$ by assuming the conditional distribution of symptoms given causes is transferable (e.g., McCormick et al.,, 2016; Byass et al.,, 2019). Such assumptions are usually violated when considering symptom distributions across deaths of different sexes and age groups (Moran et al.,, 2021) or over different populations (Li et al.,, 2024). To illustrate the conditional dependence on $D$, we use the COVID-19 dataset described in Section 2 to compute the empirical symptom prevalence and the Matthews correlation coefficient (Matthews,, 1975) of some selected symptoms across different age groups and over time. Figure 4 shows that the symptom prevalence varies across both time and age groups. For example, among COVID-19 related deaths, the proportion of observations with fever as a symptom decreases over time and as age increases. Figure 5 shows that the dependence among symptoms also varies significantly with the age and time period of the death.
@@ 3.2 Hierarchical latent class model
Figure 5: Matthews correlation coefficient (Matthews,, 1975) of a subset of symptoms among deaths related to COVID-19 among deaths in three age groups from three time periods in the Brazil VA dataset. Missing data are removed when calculating the correlation coefficient. Results for all age groups and time periods with bootstrap uncertainty intervals are reported in the Supplementary Materials.
In order to capture the heterogeneity of symptoms distributions over different sub-populations in a parsimonious way, we adopt the nested latent class model framework developed in Li et al., (2024) and Wu et al., (2024), where individual-level latent class membership indicator $Z_{i}\in\{1,2,...,K\}$ is introduced to flexibly capture the dependence of $X$. We assume the following data generating process,
@@
The introduction of the latent class nested under each cause of death leads to a flexible characterization of cause-specific dependent symptom profiles that vary across strata. That is, after integrating out the latent indicators, the conditional distribution of symptoms given a cause of death and sub-population is $p(X_{i}\mid Y_{i}=c,D_{i}=g)=\sum_{k}\lambda_{ck}^{(g)}\prod_{j}\phi_{ckj}^{x_{ij}}(1-\phi_{ckj})^{1-x_{ij}}$. When $K=1$, the model reduces to the commonly used conditional independent model in McCormick et al., (2016). In order to facilitate easier interpretations and mitigate the risk of overfitting, we fix the latent response probabilities $\bm{\phi}$ to be invariant across strata and let the sub-populations mix differently over the latent symptom profiles. With sufficiently many latent classes $K$, this representation is flexible enough to capture any multivariate discrete distribution (Dunson and Xing,, 2009).
More generally, $K$ can be treated as random and inferred from data, as is usually carried out in the Bayesian nonparametric literature (Hjort et al.,, 2010; Hannah et al.,, 2011). In this work, we fix $K=10$ to simplify the computation. Sensitivity analysis for different choices of $K$ can be found in the Supplementary Materials. For the analysis in this paper, where there are only $p=14$ symptoms, $K=10$ is sufficiently large, and posterior inference of the CSMF does not change much when $K$ is further increased.
The prior of $\pi^{(g)}$ needs to be designed for the specific choice of stratification variable $D$. We discuss our structured prior for $\pi^{(g)}$ in the next subsection. For the rest of the latent parameters, we put stick-breaking priors on $\bm{\lambda}_{c}^{(g)}$ and conjugate Beta priors on $\bm{\phi}$, i.e.,